OpenStudy (anonymous):

Find f-1(x) ***the -1 is an inverse function, not negative 1*** Sketch the graphs of y=f(x) and y=f-1(x) Equation is: f(x)=2x+3

OpenStudy (dumbcow):

y = 2x +3 what is slope? what is y-intercept?

OpenStudy (anonymous):

you can find the slope of the line by taking the derivitive of your function, f'(x)=2 but with out any known points i don't think there is a way to plug your findings back into a graphable line equation. unless you looking for a straight line that goes across y=2x, something to think about i guess. hope this helps a bit

OpenStudy (amistre64):

the slope of a straight line is its slope :) 2

OpenStudy (amistre64):

the inverse of a function is when you transpose x and y then solve for y

OpenStudy (anonymous):

yea i messed up putting the x into my y=2 equation

OpenStudy (dumbcow):

ok y-intercept is 3 so go to graph and plot point on (0,3) slope is 2 (rise/run) so from (0,3) go up 2 and over 1(2,4) repeat to get more points do the same thing for inverse funtion y=x/2 - 3/2

OpenStudy (amistre64):

f(x) = y = 2x+3 f^-1(x) = x = 2y+3 f^-1(x) = y = (x-3)/2

OpenStudy (dumbcow):

y=x/2 - 3/2 start at (0,-3/2) go up 1 over 2 (2,-1/2)